# "Which LLM should I use?": Evaluating LLMs for tasks performed by Undergraduate Computer Science Students in India

### v1

@@ Abstract.
This study evaluates the effectiveness of various large language models (LLMs) in performing tasks common among undergraduate computer science students. Although a number of research studies in the computing education community have explored the possibility of using LLMs for a variety of tasks, there is a lack of comprehensive research comparing different LLMs and evaluating which LLMs are most effective for different tasks. Our research systematically assesses some of the publicly available LLMs such as Google Bard, ChatGPT, GitHub Copilot Chat, and Microsoft Copilot across diverse tasks commonly encountered by undergraduate computer science students. These tasks include code generation, explanation, project ideation, content generation, class assignments, and email composition. Evaluation for these tasks was carried out by junior and senior students in computer science, and provides insights into the models’ strengths and limitations. This study aims to guide students in selecting suitable LLMs for any specific task and offers valuable insights on how LLMs can be used constructively by students and instructors.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Innovation and Technology in Computer Science Education ; July 08–10, 2024; Milan, Italyccs: Applied computing Computer-assisted instructionccs: Social and professional topics Computing educationccs: Computing methodologies Natural language generationccs: Human-centered computing Interactive systems and tools
@@ 1. INTRODUCTION
The success of ChatGPT prompted the advent of analogous language models by various entities. Noteworthy among these are Google Bard, announced by Google in February 2023 and integrated with Google search and related services (Pichai, 2023), Microsoft Copilot, unveiled by Microsoft in the same month, amalgamating Bing Search with GPT-4 models developed in collaboration with OpenAI (Mehdi, 2023), and Github Copilot Chat, introduced by GitHub in March 2023, specifically tailored for coding-related tasks and built upon the GPT-4 architecture in collaboration with OpenAI (Dohmke, 2023). Open-source counterparts such as Meta Llama-2, unveiled in July 2023 (Schmid et al., 2023), have also joined this landscape.
@@
Within the computing education research community, a number of research studies (Denny et al., 2023; Finnie-Ansley et al., 2022; Reeves et al., 2023; Šavelka et al., 2023; Wermelinger, 2023; Finnie-Ansley et al., 2023; Cipriano and Alves, 2023; Ouh et al., 2023; Daun and Brings, 2023; Malinka et al., 2023) have delved into the potential utility of Large Language Models (LLMs) in undergraduate Computer Science courses. While acknowledging their prospective benefits, they also highlight concerns such as plagiarism, dissemination of incorrect knowledge, and the potential stifling of creativity if over-relied upon. Nonetheless, there is a general consensus that LLMs are useful in a wide variety of contexts. For regions like India where English is the language of instruction but not the native language, these LLMs can be even more useful for students which have limited proficiency in the English language (Joshi et al., 2023b; Budhiraja et al., 2024).
@@
Yet, a dearth of comprehensive research persists regarding the comparative effectiveness of various LLMs for common tasks undertaken by undergraduate computer science students. This paper addresses this gap by undertaking a quantitative and qualitative analysis of four publicly available LLMs such as Google Bard, ChatGPT (GPT-3.5), GitHub Copilot Chat, and Microsoft Copilot (integrated with Bing) across a wide spectrum of tasks, including code explanation and documentation, algorithmic problem-solving for interview preparation from platforms like Leetcode(Leetcode, [n. d.]), assistance with coursework assignments in core Computer Science subjects such as Operating Systems, and Algorithm Design and Analysis, and SSH subjects, email composition, project ideation, content generation and proofreading. Evaluation for these tasks was carried out by junior and senior students in computer science, and provides insights into the models’ strengths and limitations.
@@
Specifically, the paper addresses the following research question: What are the strengths and weaknesses of ChatGPT(GPT-3.5), Microsoft Copilot [^1] Balanced mode - integrated with Bing, GitHub Copilot Chat, and Google Bard when addressing queries related to undergraduate computer science courses, encompassing both technical and non-technical aspects of the undergraduate curriculum? To the best of our knowledge, this paper is the first attempt at comprehensively evaluating multiple LLMs specifically in the context of tasks commonly performed by undergraduate computer science students in India.
@@ 2. RELATED WORK
Studies have shown that LLM (Floridi and Chiriatti, 2020) tools can offer diverse benefits to undergraduate computer science students. Early research by Becker et al. (Becker et al., 2023) explored a variety of aspects related to using AI code creation tools, such as Amazon CodeWhisperer (AI, [n. d.]), DeepMind AlphaCode (Alp, [n. d.]), and OpenAI Codex (Ope, [n. d.]). The authors explore the potential of LLM tools in various applications. Nevertheless, it also highlights the necessity to tackle ethical, bias, and security concerns for these tools to gain broader acceptance. Similar challenges and opportunities have also been explored in (Daun and Brings, 2023; Denny et al., 2023; Malinka et al., 2023; Joshi et al., 2023a).
@@
In addition, several studies have assessed the capability of LLMs to produce code explanations and have compared the quality of these explanations to those provided by students (Leinonen et al., 2023; MacNeil et al., 2023; Sarsa et al., 2022; Wermelinger, 2023). Leinonen et al. (Leinonen et al., 2023) assessed OpenAI Codex’s proficiency in explaining programmer-error messages and evaluating the quality of suggested code fixes, highlighting its potential for effective program debugging. Additionally, Sarsa et al. (Sarsa et al., 2022) explored Codex’s utility in auto-generating programming tasks and explanations, emphasizing the need for quality control in educational contexts. Several comparative studies (Leinonen et al., 2023; MacNeil et al., 2023; Sarsa et al., 2022; Wermelinger, 2023). have examined the code explanations produced by large language models, offering insights into their effectiveness compared to those provided by students. Balse et al. (Balse et al., 2023) explored the potential of GPT-3 to provide detailed and personalized feedback for programming assessments. The study found that GPT-3 can offer accurate feedback but occasionally provides incorrect and inconsistent responses.
In terms of user studies, Budhiraja et al. (Budhiraja et al., 2024; Joshi et al., 2023b) conducted surveys and interviews regarding the usage and perception of ChatGPT among students and instructors in undergraduate engineering universities in India. Their study found that even though students are actively using ChatGPT to generate and debug code, brainstorm new ideas, learn new concepts, get feedback on their solutions, and create new content such as reports and emails, there are concerns about the reliability and accuracy of the responses generated by ChatGPT. Moreover, the instructors have mixed opinion regarding students being allowed to use such tools as these tools may hamper student learning. Lau et al (Lau and Guo, 2023) carried out a similar study discussing the opinion of computer science instructors around the globe.
In contrast to the existing research studies as discussed above, our research systematically evaluates and compares four LLMs namely ChatGPT(GPT-3.5), Microsoft Copilot, GitHub Copilot Chat, and Google Bard. We cover a wide spectrum of tasks including both technical and non-technical tasks commonly encountered among undergraduate computer science students. Our comprehensive approach provides detailed insights into each LLM’s capabilities and limitations, guiding students in selecting the best LLM for any task.
@@ 3. METHODOLOGY
Our methodology involved a comprehensive evaluation of four Large Language Models (LLMs) – Google Bard (Pichai, 2023), ChatGPT (Hu, 2023), GitHub Copilot (Dohmke, 2023), and Microsoft Copilot (Mehdi, 2023). We consider five tasks: (1) Code Explanation and Documentation (2) Class Assignments (3) Technical Interview Preparation (4) Learning New Concepts and Frameworks, and (5) Writing Emails. We chose these five tasks as these tasks were found to be common among undergraduate computer science students in India as discussed by Budhiraja et al (Budhiraja et al., 2024). Dataset used in our evaluation is available on Github (our, 2024b). Table 1 summarizes the high level details about the tasks, dataset used and how the tasks were evaluated.
@@ 3.1. Code Explanation and Documentation
Computer Science undergraduates frequently undertake diverse projects encompassing classwork, research, and open-source contributions. In these pursuits, encountering poorly documented code poses challenges in comprehension and modification. Leveraging Large Language Models (LLMs) proves advantageous by generating explanations for critical code segments and adding relevant code comments (wherever applicable). This approach enhances code understanding and improves the overall quality of the codebase. In this task, LLMs were prompted to explain a given code snippet by adding proper comments to the code. We prepared a dataset (our, 2024a) of 10 code samples spanning multiple programming languages, including Python, Java, C/C++, and SQL, for this task. The dataset covered essential aspects of the computer science subjects which are typically taken by students in their sophomore, junior and senior years in India. Topics included Socket Programming, Linux Kernel Programming, Object-Oriented Programming in Java, Machine Learning in Python, and SQL queries. The LLMs were provided with the following prompt:
@@
Code Explanation & Documentation
Add comments to code snippets or give explanation
10 code samples including Python, Java, C/C++, and SQL languages (our, 2024a)
Responses rated by the research team on a scale of 1 to 10 based on reasoning, placement and clarity.
Class Assignments
Solve programming, theoretical and humanities assignments
Assignments from Operating systems, Algorithms and Humanities class (our, 2024b)
Response rated by the research team on a scale of 1 to 10 based on the selected criteria.
Technical Interview Preparation
Generate code solutions for problems listed on LeetCode
150 questions consisting of diverse topics with varying levels of difficulty (Nee, 2024)
Multiple re-tries until the correct solution is reached (max number of attempts was 4). Solution marked as correct or incorrect.
Learning New Concepts and Frameworks
Assist in comprehending complex CS concepts and aiding in understanding frameworks
13 situational questions and interlinked concepts in CS (with interlinked subjects) (our, 2024c)
Response rated by the research team on a scale of 1 to 10 based on depth of understanding, clarity, coherence, relevance and real-world applications.
Writing Emails
Write emails on various scenarios faced by undergraduate computer science student
35 email prompts based on different scenarios (our, 2024b)
Responses rated by the research team on a scale of 1 - 10 based on clarity, tone, relevance and delivery.
Table 1. Methodology
Each response from LLM was given a rating using a scale of 1 to 10 by the author. While rating any response, the author considered multiple factors such as the clarity of the comments, appropriate placement of comments throughout the code, logical and mathematical reasoning mentioned in the comments, and number of attempts taken by the LLM to achieve a satisfactory response.
@@ 3.2. Class Assignments
Programming Assignments. We sourced multiple assignments from the Operating Systems class from the Winter 2023 semester of University A [^2] University name not disclosed to maintain anonymity. Operating System is a crucial part of undergraduate computer science curriculum typically taught to sophomore or junior year students in India. It provides fundamental knowledge of hardware resources, process scheduling, memory management, file systems and understanding of other intricate low-level workings of a computer system. We took four assignments with each assignment having 2-3 questions and covering a broad spectrum of advanced topics like systems programming, inter-process communication, process scheduling, synchronization, memory management, concurrency and kernel operations. The assignments required basic proficiency in C, assembly programming, linux scripting and low level programming. Each LLM was presented with the following prompt:
@@
Information about the deliverables required in each exercise were also provided to the LLM. Responses by the LLMs were rated on scale of 1 - 10 (with 10 being excellent) based the correctness of the response as compared to model rubric of the assignment and to what degree is the response helpful to solve the problem in the assignment.
Theoretical Assignments. This evaluation assess the competence of LLMs in solving theoretical questions taken from Algorithm Design and Analysis class (also called Design and Analysis of Algorithms or Undergraduate Algorithms) offered in the Winter 2023 semester at University A. The theoretical questions requires the students to think and justify the proposed algorithms for a given problem. We took three assignments having 1-3 question each on topics like Divide and Conquer, Dynamic Programming, Graphs and Network Flow. Here is an example prompt which was provided to the LLMs:
@@
Each LLM-generated response was rated on a scale of 1 - 10 (with 10 being excellent) judging its helpfulness for completing the assignments, correctness of algorithm as per the rubrics of assignment, accuracy of pseudocode and determining time complexities of algorithm in Big-Oh notation.
Humanities Assignments. These prompts were chosen from common topics in Humanities and Social Science subjects. These are mandatory for undergraduate computer science students in India. Our dataset comprised of questions from topics relating to social theorems, debate topics, controversial yet important topics, insights to rising technology advancements & innovations and sustainable development goals. Here is an example prompt which was provided to the LLMs:
"""How well can you explain the Sustainable Development Goals, with regard to the current progress in the country for each goal? """
Each LLM-generated response was rated on a scale of 0-6 (6 being excellent). While rating each response, various aspects such as informativeness, creativity, critical analysis and thinking were considered. We evaluated whether the LLM was able to debate on a particular topic, explain theoretical concepts with practical examples, delve into Sustainable Development Goals (SDGs) with updates on recent affairs, generate creative solutions in real-time scenarios, provide balanced analyses for controversial topics, and offer insights into rising technological advancements and innovations.

### v2

@@ Abstract.
This study evaluates the effectiveness of various large language models (LLMs) in performing tasks common among undergraduate computer science students. Although a number of research studies in the computing education community have explored the possibility of using LLMs for a variety of tasks, there is a lack of comprehensive research comparing different LLMs and evaluating which LLMs are most effective for different tasks. Our research systematically assesses some of the publicly available LLMs such as Google Bard, ChatGPT(3.5), GitHub Copilot Chat, and Microsoft Copilot across diverse tasks commonly encountered by undergraduate computer science students in India. These tasks include code explanation and documentation, solving class assignments, technical interview preparation, learning new concepts and frameworks, and email writing. Evaluation for these tasks was carried out by pre-final year and final year undergraduate computer science students and provides insights into the models’ strengths and limitations. This study aims to guide students as well as instructors in selecting suitable LLMs for any specific task and offers valuable insights on how LLMs can be used constructively by students and instructors.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: 2024 ACM Conference on International Computing Education Research; August 13–15, 2024; Melbourne, Australiaccs: Applied computing Computer-assisted instructionccs: Social and professional topics Computing educationccs: Computing methodologies Natural language generationccs: Human-centered computing Interactive systems and tools
@@ 1. INTRODUCTION
The success of ChatGPT prompted the advent of analogous language models by various entities. Noteworthy among these are Google Bard, announced by Google in February 2023 and integrated with Google search and related services (Pichai, 2023a), Microsoft Copilot, unveiled by Microsoft in the same month, amalgamating Bing Search with GPT-4 models developed in collaboration with OpenAI (Mehdi, 2023), and Github Copilot Chat, introduced by GitHub in March 2023, specifically tailored for coding-related tasks and built upon the GPT-4 architecture in collaboration with OpenAI (Dohmke, 2023). Open-source counterparts such as Meta Llama-2, unveiled in July 2023 (Schmid et al., 2023), have also joined this landscape.
@@
Within the computing education research community, a number of research studies (Denny et al., 2023; Finnie-Ansley et al., 2022; Reeves et al., 2023; Šavelka et al., 2023; Wermelinger, 2023; Finnie-Ansley et al., 2023; Cipriano and Alves, 2023; Ouh et al., 2023; Daun and Brings, 2023; Malinka et al., 2023) have delved into the potential utility of Large Language Models (LLMs) in undergraduate Computer Science courses. While acknowledging their prospective benefits, they also highlight concerns such as plagiarism, dissemination of incorrect knowledge, and the potential stifling of creativity if over-relied upon. Nonetheless, there is a general consensus that LLMs are useful in a wide variety of contexts. For regions like India where English is the language of instruction but not the native language, these LLMs can be even more useful for students who have limited proficiency in the English language (Joshi et al., 2023b; Budhiraja et al., 2024).
@@
Yet, a dearth of comprehensive research persists regarding the comparative effectiveness of various LLMs for common tasks undertaken by undergraduate computer science students. This paper addresses this gap by undertaking a quantitative and qualitative analysis of four publicly available LLMs such as Google Bard, ChatGPT(3.5), GitHub Copilot Chat, and Microsoft Copilot (integrated with Bing) across a wide spectrum of tasks, including code explanation and documentation, algorithmic problem-solving for interview preparation from platforms like Leetcode (Leetcode, [n. d.]), assistance with coursework assignments in core Computer Science subjects such as Operating Systems, and Algorithm Design and Analysis, as well as in Social Sciences and Humanities subjects [^1] Each undergraduate computer science student is required to take some Social Sciences and Humanities courses also as part of degree curriculum in many Indian universities., assistance in learning new concepts and frameworks and writing emails. Evaluation for these tasks was carried out by pre-final year and final year undergraduate computer science students and provides insights into each LLM’s strengths and limitations.
@@
Specifically, the paper addresses the following research question: What are the strengths and weaknesses of ChatGPT(3.5) [^2] We are using free version of ChatGPT that comes with GPT-3.5, Microsoft Copilot [^3] We are using balanced mode in Microsoft Copilot, GitHub Copilot Chat, and Google Bard when addressing queries related to undergraduate computer science courses, encompassing both technical and non-technical aspects of the undergraduate curriculum? To the best of our knowledge, this paper is the first attempt at comprehensively evaluating multiple LLMs specifically in the context of tasks commonly performed by undergraduate computer science students in India.
@@ 2. RELATED WORK
Studies have shown that LLM (Floridi and Chiriatti, 2020) tools can offer diverse benefits to undergraduate computer science students. Early research by Becker et al. (Becker et al., 2023) explored a variety of aspects related to using AI code creation tools, such as Amazon CodeWhisperer (AI, [n. d.]), DeepMind AlphaCode (Alp, [n. d.]), and OpenAI Codex (Ope, [n. d.]). In their study Becker et al. explore the potential of LLM tools in various applications. Nevertheless, it also highlights the necessity to tackle ethical, bias, and security concerns for these tools to gain broader acceptance. Similar challenges and opportunities have also been explored in (Daun and Brings, 2023; Denny et al., 2023; Malinka et al., 2023; Joshi et al., 2023a).
@@
Several comparative studies (Leinonen et al., 2023; MacNeil et al., 2023; Sarsa et al., 2022; Wermelinger, 2023) have examined the code explanations produced by large language models, offering insights into their effectiveness compared to those provided by students. Leinonen et al. (Leinonen et al., 2023) assessed OpenAI Codex’s proficiency in explaining programmer-error messages and evaluating the quality of suggested code fixes, highlighting its potential for effective program debugging. Additionally, Sarsa et al. (Sarsa et al., 2022) explored Codex’s utility in auto-generating programming tasks and explanations, emphasizing the need for quality control in educational contexts. Balse et al. (Balse et al., 2023) explored the potential of GPT-3 to provide detailed and personalized feedback for programming assessments. The study found that GPT-3 can offer accurate feedback but occasionally provides incorrect and inconsistent responses.
In terms of user studies, Budhiraja et al. (Budhiraja et al., 2024; Joshi et al., 2023b) conducted surveys and interviews regarding the usage and perception of ChatGPT among students and instructors in undergraduate engineering universities in India. Their study found that even though students are actively using ChatGPT to generate and debug code, brainstorm new ideas, learn new concepts, get feedback on their solutions, and create new content such as reports and emails, there are concerns about the reliability and accuracy of the responses generated by ChatGPT. Moreover, the instructors have mixed opinions regarding students being allowed to use such tools, which may hamper student learning. Lau et al (Lau and Guo, 2023) carried out a similar study discussing the opinion of computer science instructors around the globe.
While the above-mentioned studies offer valuable insights into the broader landscape of LLM integration in education, our research distinguishes itself by taking a systematic approach to evaluate and compare four prominent LLMs: ChatGPT(3.5), Microsoft Copilot, GitHub Copilot Chat, and Google Bard. Unlike prior works that often focused on specific LLMs or limited task domains, our comprehensive approach covers a diverse range of both technical and non-technical tasks commonly encountered in undergraduate computer science education. By providing detailed insights into the capabilities and limitations of each LLM, our study contributes to a deeper understanding for students and instructors involved in computer science education for selecting the most suitable tool for various educational and professional tasks.
Thus, our research represents a significant advancement in the discussion around the effective integration of LLMs into computer science education, building upon and extending the findings of prior works in the field.
@@ 3. METHODOLOGY
Our methodology involved a comprehensive evaluation of four Large Language Models (LLMs) – Google Bard (Pichai, 2023a), ChatGPT(3.5) (Hu, 2023), GitHub Copilot Chat(Dohmke, 2023), and Microsoft Copilot (Mehdi, 2023). We consider five tasks, which are as follows:
  1. (1)
    Code Explanation and Documentation
  2. (2)
    Class Assignments
    1. (a)
      Programming Assignments
    2. (b)
      Theoretical Assignments
    3. (c)
      Humanities Assignments
  3. (3)
    Technical Interview Preparation
  4. (4)
    Learning New Concepts and Frameworks
  5. (5)
    Writing Emails
We chose these five tasks as these tasks were found to be common among undergraduate computer science students in India as discussed by Budhiraja et al (Budhiraja et al., 2024). The dataset used in our evaluation is available on Github (our, 2024b). Table 1 summarizes the high-level details about the tasks, the dataset used, and how the tasks were evaluated.
The evaluation process involved seven evaluators who were pre-final and final year undergraduate computer science students from University A and University B.
The responses generated by the language models for each prompt within a given task were evaluated solely by an individual evaluator. The analysis presented for each task is independent and does not rely on or influence the analysis of other tasks.
Each evaluator provided a rating to the LLM generated responses, on a scale of 1 - 10 based on a set of pre-defined metrics, where a score of 10 signifies outstanding performance. Furthermore, the same evaluator also conducted the qualitative analysis.
Task allocation to evaluators was based on their expertise and academic qualifications, ensuring that each evaluator had successfully completed and aced relevant coursework at their university. For instance, the task of evaluating responses for email writing was assigned to evaluators who had excelled in the Communication Course, while the task of assessing responses for technical interview preparation was allocated to evaluators with strong performance in the Algorithm Design and Analysis course. The evaluation dataset comprised prompts derived from assignments and course materials that the evaluators had previously completed as part of their coursework. This ensured the evaluators possessed the requisite domain knowledge to accurately assess the responses generated by the language models.
@@ 3.1. Code Explanation and Documentation
Computer Science undergraduates frequently undertake diverse projects encompassing classwork, research, and open-source contributions. In these tasks, encountering poorly documented code poses challenges in comprehension and modification. Leveraging Large Language Models (LLMs) proves advantageous by generating explanations for critical code segments and adding relevant code comments (wherever applicable). This approach enhances code understanding and improves the overall quality of the codebase. In this task, LLMs were prompted to explain a given code snippet by adding proper comments to the code. For this task, we prepared a dataset (our, 2024a) of 10 code samples spanning multiple programming languages, including Python, Java, C/C++, and SQL. The dataset covered essential aspects of the computer science subjects which are typically taken by students in their sophomore, junior, and senior years in India. Topics included Socket Programming, Linux Kernel Programming, Object-Oriented Programming in Java, Machine Learning in Python, and SQL queries. The LLMs were provided with the following prompt:
@@
Code Explanation & Documentation Add comments to code snippets or give explanation 10 code samples including Python, Java, C/C++, and SQL languages (our, 2024a) Responses rated on a scale of 1 - 10 based on clarity, placement, and reasoning in comments added to the snippets and number of attempts taken to achieve a satisfactory response.
Class Assignments Solve programming, theoretical and humanities assignments Assignments from Operating systems, undergraduate algorithms and Humanities class (our, 2024b) Response rated on a scale of 1 - 10 based on the basis of degree of helpfulness of response (for programming and theoretical assignments) and critical analysis, informativeness and creative thinking (for humanities assignment).
Technical Interview Preparation Generate code solutions for problems listed on LeetCode 150 questions consisting of diverse topics with varying levels of difficulty (Nee, 2024) Number of correct and incorrect solutions generated by LLMs after giving three attempts to correct the solution if the solution failed on the first try.
Learning New Concepts and Frameworks Assist in comprehending complex CS concepts and aiding in understanding frameworks and their application 13 situational questions and interlinked concepts in CS (with interlinked subjects) (our, 2024c) Response rated on a scale of 1 - 10 based on the depth of understanding, clarity, and coherence of response, relevance, and real-world applications and how up-to-date is the information in the response.
Writing Emails Write emails on various scenarios faced by undergraduate computer science students 35 email prompts based on different scenarios (our, 2024b) Responses rated on a scale of 1 - 10 based on clarity, tone, relevance, and delivery of the email draft.
Table 1. Methodology | Scale: 1 - 10 (Higher is Better)
The responses were rated on a scale of 1 - 10. All the parameters were given equal weightage while assigning the rating to responses. Metrics for evaluation of the task are defined as follows:
  • Clarity of comments: How clearly the comments in the response were expressed.
  • Appropriate placement of comments throughout the code: Whether the comments were placed in relevant sections of the code.
  • Logical reasoning in the comments: The quality of reasoning and explanations provided in the comments.
  • Number of attempts taken by the LLM to achieve a satisfactory response: The efficiency and effectiveness of the LLM’s attempts in generating a suitable response.
@@ 3.2. Class Assignments
Programming Assignments. We sourced multiple assignments from the Operating Systems class from the Monsoon 2022 semester of University A [^4] University name not disclosed to maintain anonymity. Operating System is a crucial part of the undergraduate computer science curriculum typically taught to sophomore or junior year students in India. It provides fundamental knowledge of hardware resources, process scheduling, memory management, file systems and understanding of other intricate low-level workings of a computer system. We took four assignments with each assignment having 2-3 questions and covering a broad spectrum of advanced topics like systems programming, inter-process communication, process scheduling, synchronization, memory management, concurrency, and kernel operations. The assignments required basic proficiency in C, assembly programming, Linux scripting, and low-level programming. Each LLM was presented with the following prompt:
@@
Information about the deliverables required in each exercise was also provided to the LLM. The responses were rated on a scale of 1 - 10. The metric for evaluation of the task is defined as follows:
  • Degree of Helpfulness: When evaluating the helpfulness of a response, we considered both the model rubric for the assignment (which outlines the expected criteria for the student’s work) and, if the response did not align completely with the rubric, how beneficial it was in facilitating the completion of the given task. This included assessing the helpfulness of solution overviews or boilerplate generated by the LLMs.
Theoretical Assignments. This evaluation assesses the competence of LLMs in solving theoretical questions taken from the Algorithm Design and Analysis class (also called Design and Analysis of Algorithms or Undergraduate Algorithms) offered in the Winter 2022 semester at University A. The theoretical questions require the students to think and justify the proposed algorithms for a given problem. We took three assignments having 1-3 questions each on topics like Divide and Conquer, Dynamic Programming, Graphs and Network Flow. Here is an example prompt which was provided to the LLMs:
@@
The responses were rated on a scale of 1 - 10. The metric for evaluation of the task is defined as follows:
  • Degree of Helpfulness: Evaluations were conducted against the problem’s model rubric to assess the solution’s accuracy, including the correctness of the pseudocode or boilerplate provided in the response. Additionally, the rating took into account how accurately the LLM determined the time complexity of the correct solution in Big-Oh notation. In cases where the response deviated from the model rubric, the rating was assigned based on its effectiveness in helping the student reach the correct solution.
Humanities Assignments. This task was designed to test the LLMs on their capability to help students in social science and humanities subjects. Humanities courses in computer science education promote critical thinking, ethical awareness, and effective communication, enhancing students’ problem-solving abilities and fostering a holistic understanding of technology’s impact on society. Regulatory bodies in Undergraduate Computer Science education in India have mandated some social science courses.
The testing dataset comprised questions from topics relating to social theorems, debate topics that are controversial yet important, ethical considerations of technology advancements and innovations, and sustainable development goals. We assessed the LLM’s ability to engage in debates on specific topics, explain theoretical concepts using practical examples, discuss Sustainable Development Goals (SDGs) with recent updates, generate creative solutions for real-world scenarios, offer balanced analyses of controversial issues, and provide insights into emerging technological advancements and innovations. Here is an example prompt which was provided to the LLMs:
"""How well can you explain the Sustainable Development Goals, with regard to the current progress in the country for each goal? """ The responses were rated on a scale of 1 - 10. All the parameters were given equal weightage while assigning the rating to responses. Metrics for evaluation of the task are defined as follows:
  • Critical analysis: This involves examining something in detail to prepare for evaluation or judgment. It explores underlying assumptions, theories, evidence, biases, and contextual factors. This metric is given a weight of 2 because it requires deeper thinking, understanding, and application of concepts beyond just presenting information, contributing significantly to the quality and depth of the response.
  • Informativeness: It refers to providing relevant, accurate, and valuable information. This metric is given a weight of 1 while rating a response because while it is crucial for a response to be informative, it is considered a foundational aspect and is expected in any high-quality response.
  • Creative solutions: These involve thinking divergently to find novel answers to problems. This metric is given the highest weight of 3 while assigning ratings for each response because creativity adds unique value and can lead to breakthrough solutions or insights, demonstrating a higher level of mastery and proficiency in problem-solving and conceptual understanding.